OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (anonymous):

Can you help? @pooja195

OpenStudy (anonymous):

1st term: f(1) easy 2nd term: f(2) easy 3rd term: f(3) = f(2) + f(1) = ? 4th term: f(4) = f(3) + f(2) = ? you get the idea?

OpenStudy (anonymous):

so the 3rd term = f(1) right?

OpenStudy (anonymous):

so my answer is B? @Lindsayb12

OpenStudy (solomonzelman):

yup

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

ok thanks!!